OpenStudy (anonymous):

if one pulley is 8 inches bigger in diameter than another and they are connected by a belt, how big are the pulleys if the smaller pulley turns at 120 rpm and the larger turns at 40rpm..? Someone plz help

OpenStudy (anonymous):

if theyre connected by a belt then r1ω1 = r2ω2 r2 = r1 + 8 120 x r1 = 40 x (r1+8)\ r1=4, r2=12

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

no prob